QUESTION 2 the joule is equivalent to which of the units below? O m/N kg·m/s2 N/m kg.m2/s2
Answer #1

Joule is the unit of energy:

Now Energy = Force x displacement

Unit of force = Newton(N) = kg m/s2

Unit of displacement = meter(m)

Hence energy = kg m/s2 x m = kg-m2/s2

Hence Joule = kg-m2/s2